Toss two fair coins and record the outcome. Find the probability of observing exactly one head in the two tosses.
Answer #1

Two fair coins are tossed.

The all possible outcome is {HH,HT,TH,TT}

That is 2^{}^{}^2 =4

We want the probability of exactly one head in two tosses.

So there is only two such outcome for this event ie {HT,TH}.

P(exactly one head) =2/4 = 1/2